Dry hydrogen by pellet of mercury in a uniform capillary tube closed at one end. If the length of the column of hydrogen at 27 is 1.0 m, at what temperature will the length be 1.20 m
1
Expert's answer
2020-06-02T17:54:43-0400

Use the constant pressure law for the two states of the unchanged mass of gas:


"\frac{V_1}{T_1}=\frac{V_2}{T_2},\\
\space\\
T_2=\frac{V_2}{V_1}T_1=\frac{1.2\cdot A}{1\cdot A}(273+27)=360\text{ K}."

This temperature is equivalent to 87°C.